Gallatin Carpet Cleaning is a small, family-owned business operating out of Bozeman, Montana. For its services, the company hResource consumption is distributed across the activities as follows: Job Support 0% Other 15% Distribution of Resource ConsuComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Prepare the fiComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the acComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 The company reComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 The revenue fr
Answer #1
1 Cleaning Carpets Travel to Jobs Job support Other Total
Wages                  99,280             16,320                 -       20,400 136,000
Cleaning Supplies                  35,000                     -                   -              -       35,000
Cleaning Equipment depreciation                    4,620                     -                   -         2,380       7,000
Vehicle Expense                          -               27,060                 -         5,940     33,000
Office Expense                          -                       -            44,160     24,840     69,000
President's Compensation                          -                       -            26,560     56,440     83,000
Total Cost                138,900             43,380          70,720 110,000 363,000
2 Activity Rates for the Activity cost Pools
Activity Cost Pool Expected Overhead Costs Expected Activity Cost Activity Rate
Cleaning Carpets                               138,900                             8,500             16.34 Per hundred Sq. feet
Travel to jobs                                 43,380                         433,000               0.10 Per Mile
Job support                                 70,720                             1,800             39.29 Per Job
                              253,000
3 Cost of the Job
Job Usage of Activity in the Job Activity Rate Applied Cost to Job
Cleaning Carpets                                       600             16.34                    9,804.71
Travel to jobs                                         51               0.10                           5.11
Job support                                           1             39.29                         39.29
Total cost of the Job                    9,849.10
4 Customer Margin on this Job
Revenue 132.90
Cost of the Job per sq. feet     16.42
Customer Margin 116.48

Please find below table useful to compute desired results: -

B А 1 Current Fee 2 Activity Cost Pool 3 Cleaning Carpets 4 Travel to jobs 5 Job support 6 Other 7 Wages 8 Cleaning Supplies3 H 1 1] Cleaning Carpets 2 Wages =B7*B16 Cleaning Supplies =SB8*B17 Cleaning Equipment depreciation =SB9*B18 Vehicle Expense

End results would be as follows: -

EF А 1 Current Fee 2 Activity Cost Pool 3 Cleaning Carpets 4. Travel to jobs 5 Job support 6 Other 7 Wages 8 Cleaning SupplieCleaning Carpets Travel to Jobs Job support Other 16,320 20.400 99,280 35,000 4.620 Wages Cleaning Supplies Cleaning Equipmen